OpenStudy (anonymous):

double integral of x^3cos(xy)dydx from o to pi/4 and sinx to cosx

OpenStudy (anonymous):

\[\pi/4\] ?

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi/4}\int\limits_{sinx}^{cosx} x^3 \cos(xy) dydx\]

OpenStudy (anonymous):

lol yeah, that's why i'm asking for help :o

OpenStudy (anonymous):

i have no idea how to integrate that partially lol

OpenStudy (wasiqss):

firstly integrate with respect to x, keeping y constant

OpenStudy (wasiqss):

will be done by by parts

OpenStudy (anonymous):

why am i integrating with respect to x first, if its dydx? o.O

OpenStudy (wasiqss):

thats the way partial integrals are done, first we need to do wrt to dx, and put in limits then whatever the answer we get we have to integrate wrt to y, and put in the limiits , that will be final answer

OpenStudy (anonymous):

but isn't that when its dxdy? when it's dydx you integrate wrt y first, and then x? o.O

OpenStudy (wasiqss):

ohhh i see, yeh we need to do that way, and yeh we have to interchange the limit dude

OpenStudy (anonymous):

if i integrate with respect to x first, then do i need to change the limits?

OpenStudy (wasiqss):

wait i will do simplified form for you

OpenStudy (anonymous):

so if its y = sinx then ... x = sin*-1y?!

OpenStudy (wasiqss):

\[\int\limits_{?}^{y/\sqrt{2}}\int\limits_{0}^{1/\sqrt{2}}x^3\cos(xy)dxdy\]

OpenStudy (wasiqss):

now do it..........

OpenStudy (anonymous):

wait.. how did you get that? o.O

OpenStudy (wasiqss):

btw in which year in college you are?

OpenStudy (anonymous):

senior, why?

OpenStudy (wasiqss):

well its called interchanging the limit,

OpenStudy (wasiqss):

because we need to have dx first, for simplification

OpenStudy (anonymous):

wait, why is having dx first simplified? and would you mind explaining to me how you interchanged the limit, if you don't mind?

OpenStudy (wasiqss):

because , according to question, we have to integrate wrt to y first, but you can see , for the first integration, the limits are given in terms of x, so wee need to interchange

OpenStudy (anonymous):

wait lets try this with a much simplier integration, cuz i'm getting really confused with what you mean.. so lets say the limits are the same as the problem given above, but we integrate 1 okay? \[\int\limits_{0}^{\Pi/4}\int\limits_{sinx}^{cosx}1 dydx = y |(from sinx \to cosx)\] = cosx - sinx \[=\int\limits_{0}^{\Pi/4} cosx - sinx dx\] = sinx + cosx | (from 0 to Pi/4) see what i mean? the limits, although in terms of x, are actually functions of y y = sinx, and y = cosx, that's why it's dydx i don't think we should be changing the limits =\ do you know how to partially integrate the equation in terms of y first? :o the cos(xy) is killing me lol

OpenStudy (anonymous):

Is this the question no one has answered?

OpenStudy (wasiqss):

it is in pending

OpenStudy (anonymous):

yes :(

OpenStudy (anonymous):

@KingGeorge i'm so happy you're online!!

OpenStudy (kinggeorge):

alright, integrating dydx would be easier So we have the following integral \[\large \int\limits\limits_{0}^{\pi/4}\int\limits\limits_{\sin x}^{\cos x} x^3 \cos(xy)\;\; dydx\]If we integrate with respect to y first, we get \[\large \int\limits_{0}^{\pi/4} \left(x^2 \sin(xy) |_{\sin x}^{\cos x}\right) dx\]

OpenStudy (anonymous):

wait.. seriously? you don't need to do anything with the (xy)? you just integrate the cos? .. that's it? thats what i've been struggling so much witH? lol..

OpenStudy (kinggeorge):

since you can consider x a constant, you're just integrating \(\cos(xy)\) with respect to y.

OpenStudy (anonymous):

when i'm plugging it in.. is it just x^3sin(cosxx? lol

OpenStudy (anonymous):

cosx^2?

OpenStudy (kinggeorge):

Not quite. I'm getting \[x^2\sin(x \cos(x))-x^2\sin(x \sin(x))=x^2(\sin(x \cos(x))-\sin(x \sin(x))\]

OpenStudy (anonymous):

okay so.. how the heck.. do i integrate this monstrosity? lol

OpenStudy (wasiqss):

lool

OpenStudy (kinggeorge):

I have no idea of how to integrate that monstrosity by hand.

OpenStudy (anonymous):

ytfreak is it in our book.. lol

OpenStudy (anonymous):

okay what about integrating x^3sin(x^3)?

OpenStudy (experimentx):

another montrosity http://www.wolframalpha.com/input/?i=integrate+x^3+sin%28x^3%29

OpenStudy (anonymous):

wtfreak is this. what is my teacher thinking...

OpenStudy (anonymous):

i?!? there's an i int here!? you can get an i when you integrate?! what is going on.

OpenStudy (kinggeorge):

Perhaps your teacher meant \(x^2 \sin(x^3)\)?

OpenStudy (anonymous):

but isn't that really hard to do as well? did you see what @experimentX linked? lol

OpenStudy (kinggeorge):

\(x^2 \sin(x^3)\) is actually pretty easy.

OpenStudy (kinggeorge):

\[\int\limits x^2 \sin(x^3)\;\;dx =-{1 \over 3} \cos(x^3)+C\]

OpenStudy (anonymous):

what... how did you get that? :o

OpenStudy (kinggeorge):

I've seen enough integrals to recognize that pattern, but analytically, a u-sub of \(u=x^3\) would do the trick.

OpenStudy (anonymous):

oooohhhhhhhh lol duh... -_-"